OpenStudy (anonymous):

Hi everyone, would someone be able to help me with the following: Exact Interest for Principle of $3400, Rate is 11% and Time is 176 Days. I keep getting $179.52 which is not a choice on my homework. Any help is greatly appreciated, denise

OpenStudy (anonymous):

Rate = I/PT. That I can figure but somewhere I am messing up, thanks again, denise PS Exact interest is used as 365 days, I've been doing that part right too, I think, lol!

OpenStudy (anonymous):

Oh lord, I'm too tired, I messed up the formula I see

OpenStudy (klimenkov):

11% means that someone will obtain $3400*0.11=$374 in a year. But 176 days is a \(\frac{176}{365}=0.482\) part of the year. So it will be $374*0.482=$180,268.

OpenStudy (anonymous):

What is the exact Interest?

OpenStudy (anonymous):

P X R x t

OpenStudy (klimenkov):

I think, it is. Do you have an answer to this?

OpenStudy (anonymous):

yes, one sec

OpenStudy (anonymous):

18.03 or 182.84 or 18.28 or 18.34

OpenStudy (anonymous):

I get 180.27?

OpenStudy (anonymous):

3400 X .11 X .482 ?

OpenStudy (klimenkov):

It is the first one. Using the calculator with high precision: http://www.wolframalpha.com/input/?i=3400*0.11*176%2F365

OpenStudy (anonymous):

Dangit, I knew it was going to be my calc. I can't seem to get it right, but hey, thank you much, I will try using the online calculator until I can handle this one I bought:) ty again, denise

OpenStudy (klimenkov):

You are welcome. If you have answer, don't be shy to ask them.